# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@ 2.2 The order on $X_{-1}$
Let $p\in(1,\infty)$ and let $A:L^{p}(\Omega)\supseteq\operatorname{dom}A\to L^{p}(\Omega)$ denote the same operator as in Example 2.20, but now with Dirichlet boundary conditions. Assume in addition that the bounded domain $\Omega\subseteq\mathbb{R}^{d}$ has $C^{2}$-boundary and that the coefficient matrix $a$ is symmetric and $C^{1}$ on $\overline{\Omega}$. Again, it is well-known that $A$ generates a positive and analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $L^{p}(\Omega)$.
Assume now that $p\geq 2$ and that $\alpha:=\frac{1}{2q}(1+\frac{d}{2})<1$ (note that both conditions can only be simultaneously satisfied if $d\leq 5$); here $q$ denotes the Hölder conjugate of $p$. Then $L^{p}(\Omega)_{-1,+}$ is contained in the negative Favard space $F_{-\alpha}$ and thus in a real interpolation space between $X$ and $X_{-1}$.
@@
To prove this, first note that since the coefficient matrix $a$ is assumed to be symmetric, the dual operator $T(t)^{\prime}$ acts as the continuous extension of $T(t)$ to the space $L^{q}(\Omega)$ for the Hölder conjugate $q$ of $p$. Let $u\in L^{p}(\Omega)_{+}$ denote the leading eigenfunction of $A$ and thus also of $A^{\prime}$ (so one also has $u\in L^{q}(\Omega)_{+}$) associated to the leading eigenvalue $-\lambda_{0}\in(-\infty,0)$. We show that the assumptions of Theorem 2.18(a) are satisfied for the functional $\varphi:=u\in L^{q}(\Omega)$.
@@
The smoothness assumptions on $\partial\Omega$ and on the coefficient matrix $a$ imply that the heat kernel $k_{t}:\Omega\times\Omega\to[0,\infty)$ of the semigroup operator $T(t)$ (and thus also of $T(t)^{\prime}$) satisfies the intrinsic ultracontractivity estimate
(2.1) $k_{t}(x,y)\leq ct^{-\frac{1}{2}(1+\frac{d}{2})}u(x)u(y)=ct^{-\alpha q}u(x)u(y)$
@@
for a constant $c>0$, all $t>0$, and all $x,y\in\Omega$ [14, Theorem 4.6.2] (it should be noted that there is a typo in the formula in [14, Theorem 4.6.2]: the factor $-1$ in the exponent should actually be $-\frac{1}{2}$, as the proof shows). Hence, $T(t)^{\prime}$ maps each $L^{q}(\Omega)$ into the principal ideal $(L^{q}(\Omega))_{u}$. Moreover, since $u\in\operatorname{dom}A^{\prime}$, the space $\operatorname{dom}A^{\prime}\cap(L^{q}(\Omega))_{u}$ is clearly majorizing in $(L^{q}(\Omega))_{u}$. So it only remains to show that $\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to(L^{q})_{u}}\leq\tilde{c}t^{-\alpha}$ for a constant $\tilde{c}>0$ and all $t\in(0,1]$.
@@
To this end, we first observe that the semigroup $(T(t))_{t\geq 0}$ is dominated by the semigroup generated by the same differential operator but with Neumann boundary conditions; this follows for instance from [44, Corollary 2.22]. In addition to the principal ideals $(L^{p}(\Omega))_{u}$ and $(L^{q}(\Omega))_{u}$ we will now use the weighted $L^{1}$-space $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$. This space coincides with the norm completion of $L^{p}(\Omega)$ with respect to the norm $\left\lVert f\right\rVert_{L^{1}(\mathrm{d}x)}:=\int_{\Omega}\left\lvert f\right\rvert u\;\mathrm{d}x$ and behaves, in a sense, dually to the principal ideal $(L^{q}(\Omega))_{u}$; we refer to [9, Section 2] or [13, Section 2] for a detailed discussion.
@@
Moreover, since $T(t)^{\prime}u=e^{-t\lambda_{0}}u$ for each $t$, the operator $T(t)$ extends to a bounded linear operator on $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$, which we again denote by $T(t)$ and which satisfies
@@
(an abstract version of this argument can be found in [9, Proposition 2.2]). We will now combine the previous two norm estimates for $T(t)$ with an interpolation inequality to get an estimate for $\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to L^{p}}$: for every function $f\in(L^{p}(\Omega))_{u}\subseteq L^{p}(\Omega)\subseteq L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$ we have
@@
so $\left\lVert f\right\rVert_{L^{p}}\leq\left\lVert u^{p-2}\right\rVert_{\infty}^{1/p}\left\lVert f\right\rVert_{L^{1}(u\;\mathrm{d}x)}^{1/p}\left\lVert f\right\rVert_{(L^{p})_{u}}^{1/q}$; here we have also used that $u\in L^{\infty}(\Omega)$ and the assumption $p\geq 2$, which ensures that $\left\lVert u^{p-2}\right\rVert_{\infty}<\infty$. This estimate for $\left\lVert f\right\rVert_{L^{p}}$ together with the two aforementioned norm estimates for $T(t)$ readily gives the interpolated norm estimate
$\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to(L^{p})_{u}}\leq c^{1/q}\left\lVert u^{p-2}\right\rVert_{\infty}^{1/p}t^{-\alpha}$
for all $t>0$. For every $t>0$, one has $\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to(L^{p})_{u}}=\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to(L^{q})_{u}}$. Indeed, the inequality $\leq$ is shown in an abstract setting in [13, Proposition 2.1], but one can show that even equality is true by using the duality result in [47, Exercise IV.9(a)]. So the assumption of Theorem 2.18(b) is satisfied, which gives the claimed result.
For Dirichlet boundary conditions, not all elements of $L^{p}(\Omega)_{-1,+}$ can be described as measures on $\overline{\Omega}$. This is in contrast to the situation for Neumann boundary conditions and small $p$, see Example 2.20(b). One can see this easily by taking $A$ to be the Dirichlet Laplacian on the interval $(-1,1)$. Then $A^{\prime}$ is also the Dirichlet Laplacian and the domain $\operatorname{dom}A^{\prime}$ is (no matter which value we use for $p$) a majorizing and dense subset of $C^{1}[-1,1]$. Hence, the linear functionals $f\mapsto f^{\prime}(-1)$ and $f\mapsto-f^{\prime}(1)$ are positive elements of $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}((-1,1))_{-1}$ that are not given by measures on $[-1,1]$.
3 Admissibility of observation operators
We now come to the main topic of the paper – admissibility. Throughout this section, we consider the linear time-invariant system
@@ 3 Admissibility of observation operators
where $A$ is the generator of a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on a Banach space $X$ and the observation operator $C\in\mathcal{L}(X_{1},Y)$ takes values in a Banach space $Y$. The output map is defined as
(3.1) $\Psi_{\tau}:X_{1}\to L^{1}([0,\tau],Y),\qquad x\mapsto CT(\mathord{\,\cdot\,})x$
for a fixed $\tau>0$. We say that $C$ is a $\mathrm{L}^{1}$-admissible observation operator if $\Psi_{\tau}$ has a bounded extension to $X$ for some (equivalently, all) $\tau>0$. This is of course, equivalent to the existence of $K_{\tau}>0$ such that
(3.2) $\left\lVert CT(\mathord{\,\cdot\,})x\right\rVert_{L^{1}([0,\tau],Y)}\leq K_{\tau}\left\lVert x\right\rVert,\qquad(x\in X_{1}).$
In addition, if we have $\limsup_{\tau\downarrow 0}\left\lVert\Psi_{\tau}\right\rVert_{\mathcal{L}(X,\mathrm{L^{1}}([0,\tau],Y))}=0$, then $C$ is called a zero-class $\mathrm{L^{1}}$-admissible observation operator. Most of our results in this section and the sequel are proved in the following setting:
{assumption}
@@
In all of our results, the order on $X_{1}$ is assumed to be the order which it inherits from $X$. We begin with obtaining sufficient conditions for an operator $C\in\mathcal{L}(X_{1},Y)$ to be $L^{1}$-admissible. Thereafter, in Section 3.2, we look at situations where $L^{1}$-admissibility of $C$ automatically implies zero-class $L^{1}$-admissibility.
3.1 $L^{1}$-admissibility
For positive semigroups on Banach lattices, we know from [56, Theorem 4.1.17] that positive observation operators mapping into AL-spaces are always $L^{1}$-admissible. In our first result, we generalise this to ordered Banach spaces whose norm is additive on the positive cone.
@@ 3.1 $L^{1}$-admissibility
Suppose that Assumption 3 holds. Assume further that $Y$ is an ordered Banach space whose norm is additive on the positive cone and that $C\in\mathcal{L}(X_{1},Y)$ is positive. Then $C$ is $L^{1}$-admissible.
For the proof of Proposition 3.1, we essentially adapt the argument of [56, Theorem 4.1.17] to our setting. The essential ingredient of the argument is that due to the positivity of the semigroup and the observation operator, it suffices to establish (3.2) for positive $x\in X_{1}$. When $X$ and $Y$ are both Banach lattices, this was proved in [19, Lemma 3.1]. However, the result remains true under our assumption as well and is actually a consequence of the following technical lemma.
@@
Let $X$ be an ordered Banach space with a generating cone, let $\widetilde{X}$ be densely embedded in $X$, and let $Y$ be a Banach space. Let $(F(t))_{t\geq 0}$ be a family of operators in $\mathcal{L}(\widetilde{X},Y)$ with the following properties:
@@
    If $x\in\widetilde{X}$ and $y,z\in X_{+}$ are such that $x=y-z$, then there exist sequences $(y_{n})$ and $(z_{n})$ in $\widetilde{X}_{+}$ such that $y_{n}\to y$ and $z_{n}\to z$ in $X$ and $\left\lVert F(t)(y_{n}-z_{n})\right\rVert\to\left\lVert F(t)x\right\rVert$ for every $t\geq 0$.
@@
    The mapping $t\mapsto F(\mathord{\,\cdot\,})x$ lies in $L^{1}([0,\tau],Y)$ for each $x\in\widetilde{X}$.
[…]
    There exists $\widetilde{M}>0$ such that $\int_{0}^{\tau}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq\widetilde{M}\left\lVert x\right\rVert$ for all $x\in\widetilde{X}_{+}$.
Then there exists $M>0$ such that
$\int_{0}^{\tau}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq M\left\lVert x\right\rVert$
for all $x\in\widetilde{X}$.
Proof 3.3.
Let $x\in\widetilde{X}$. By the uniform decomposition property of ordered Banach spaces with generating cones [4, Theorem 2.37], there exists $c>0$ – independent of $x$ – and $y,z\in X_{+}$ such that $x=y-z$ and $\left\lVert y\right\rVert,\left\lVert z\right\rVert\leq c\left\lVert x\right\rVert$. Let $(y_{n})$ and $(z_{n})$ be sequences as in assumption (a). Then
$\int_{0}^{\tau}\left\lVert F(t)(y_{n}-z_{n})\right\rVert\;\mathrm{d}t\leq\widetilde{M}\left(\left\lVert y_{n}\right\rVert+\left\lVert z_{n}\right\rVert\right)$
for each $n\in\mathbb{N}$. An application of the dominated convergence theorem now implies that
$\int_{0}^{\tau}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq\widetilde{M}(\left\lVert y\right\rVert+\left\lVert z\right\rVert)\leq 2c\widetilde{M}\left\lVert x\right\rVert,$
as desired.
Proof 3.4 (Proof of Proposition 3.1).
Without loss of generality, assume that $\omega_{0}(A)<0$ and let $C\in\mathcal{L}(X_{1},Y)$ be positive. Then for every positive $x\in X_{1}$, we have
[…]
here the first equality holds because the norm on the positive cone of $Y$ is additive.
On the other hand, if $x\in X_{1}$, there exist $y,z\in X_{+}$ such that $x=y-z$ because the cone of $X$ is generating. Define $y_{n}:=nR(n,A)y$ and $z_{n}:=nR(n,A)z$. Then $(y_{n})$ and $(z_{n})$ are positive sequences in $X_{1}$ converging respectively to $y$ and $z$ in $X$ such that for each $t\geq 0$, the sequence $x_{n}:=y_{n}-z_{n}$ satisfies
$\Big{\lvert}\left\lVert CT(t)x_{n}\right\rVert-\left\lVert CT(t)x\right\rVert\Big{\rvert}$ $=\Big{\lvert}\left\lVert CA^{-1}T(t)Ax_{n}\right\rVert-\left\lVert CA^{-1}T(t)Ax\right\rVert\Big{\rvert}$
$\leq\left\lVert CA^{-1}\right\rVert\left\lVert T(t)\right\rVert\left\lVert Ax_{n}-Ax\right\rVert\to 0$
as $n\to\infty$. The assertion is now an immediate consequence of Lemma 3.2.
@@
Suppose that Assumption 3 holds. Assume further that there exists an ordered Banach space $\widetilde{X}$ whose norm is additive on the cone $\widetilde{X}_{+}$ such that $C$ factorises as
@@
Proof 3.6.
By Proposition 3.1, the operator $C_{1}$ is $L^{1}$-admissible. Thus there exists $K_{\tau}>0$ such that
@@
for all $x\in X_{1}$, As a result, $C=C_{2}C_{1}$ is also $L^{1}$-admissible.
The factorisation condition in the preceding corollary might seem a bit artificial at first glance. However, we now give two corollaries which show that the condition is quite natural in some situations.
@@
Suppose that Assumption 3 holds. If $Y$ is an ordered Banach space and $C\in\mathcal{L}(X_{1},Y)$ is a positive finite-rank operator, then $C$ is $L^{1}$-admissible.
Proof 3.8.
To apply Theorem 3.5, we define $\widetilde{X}\subseteq Y$ to be the span of the range $\operatorname{Rg}C$ and we endow it with the cone $\widetilde{X}_{+}:=\widetilde{X}\cap Y_{+}$. Then $\widetilde{X}$ is a finite-dimensional ordered Banach space and hence, there exists an equivalent norm on this space that is additive on the positive cone; we refer to [4, Corollary 3.8] and [57, Theorem VII.1.1]. So if we set $C_{1}:=C$ (but with co-domain $\widetilde{X}$) and let $C_{2}$ be the canonical embedding of $\widetilde{X}$ into $Y$, Theorem 3.5 is applicable.
One can also use the factorisation assumption in Theorem 3.5 in the following way: the dual statement of this assumption can be characterised in terms of ordered boundedness, as we show in Appendix A. If $Y$ is reflexive, one can thus reformulate Theorem 3.5 in the following way.
@@
Suppose that Assumption 3 holds and assume that $Y$ is reflexive. If the dual operator $C^{\prime}:Y^{\prime}\to(X_{1})^{\prime}=(X^{\prime})_{-1}$ maps the unit ball of $Y^{\prime}$ into an ordered bounded subset of $(X^{\prime})_{-1}$, then $C$ is $L^{1}$-admissible.
Proof 3.10.
@@
$C:X_{1}\overset{C_{2}^{\prime}}{\longrightarrow}(\widetilde{X})^{\prime}\overset{C_{1}^{\prime}}{\longrightarrow}Y,$
where $C_{2}^{\prime}$ is positive because $C_{2}$ is.
Finally, since $\widetilde{X}$ has a unit, so the norm on $(\widetilde{X})^{\prime}$ is additive on its positive cone [42, Lemmata 2 and 3]. We have thus shown that all assumptions of Theorem 3.5 are fulfilled and thus $C$ is $L^{1}$-admissible.
3.2 Zero-class $L^{1}$-admissibility
While $L^{1}$-admissible observation operators need not be zero-class $L^{1}$-admissible (see Example 3.16 below), we are interested in situations when this is indeed the case. To this end, we take advantage of the nilpotency of the right-translation semigroup on $L^{1}([0,1],Y)$ to formulate a simple test for $L^{1}$-admissibility to be zero-class.

### v4

@@ 2.2. The order on $X_{-1}$
Let $p\in[2,\infty)$, let $\Omega\subseteq\mathbb{R}$ be a bounded open interval, and let $A:L^{p}(\Omega)\supseteq\operatorname{dom}A\to L^{p}(\Omega)$ denote the same operator as in Example 2.13, but now with Dirichlet boundary conditions. Assume in addition that the coefficient $a$ is $C^{1}$ on $\overline{\Omega}$. Again, it is well-known that $A$ generates a positive and analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $L^{p}(\Omega)$. Let $q\in(1,2]$ be the Hölder conjugate of $p$ and suppose $\alpha:=\frac{3}{2q}<1$ (all of the following arguments would also work on smooth $d$-dimensional domains if the coefficient $a$ is symmetric and $\alpha:=\frac{1}{q}(1+\frac{d}{2})<1$; but this inequality in conjunction with $q\leq 2$ is possible only if $d=1$).
@@
We show that $L^{p}(\Omega)_{-1,+}$ is contained in the Favard space $F_{-\alpha}$. We begin by noting that since the coefficient matrix $a$ is assumed to be symmetric, the dual operator $T(t)^{\prime}$ acts as the continuous extension of $T(t)$ to $L^{q}(\Omega)$. Let $u\in L^{p}(\Omega)_{+}$ denote the leading eigenfunction of $A$ and thus also of $A^{\prime}$ (in particular, $\varphi:=u\in L^{q}(\Omega)_{+}$) associated to the leading eigenvalue $-\lambda_{0}\in(-\infty,0)$. We show that the assumptions of Theorem 2.12(a) are satisfied.
@@
The smoothness assumption on the coefficient $a$ implies that the heat kernel $k_{t}:\Omega\times\Omega\to[0,\infty)$ of the semigroup operator $T(t)$ (and thus also of $T(t)^{\prime}$) satisfies the intrinsic ultracontractivity estimate
$k_{t}(x,y)\leq ct^{-\frac{3}{2}}u(x)u(y)=ct^{-\alpha q}u(x)u(y)$ (2.1)
@@
for a constant $c>0$, all $t>0$, and all $x,y\in\Omega$ [17, Theorem 4.6.2]. Hence, $T(t)^{\prime}$ maps each $L^{q}(\Omega)$ into the principal ideal $(L^{q}(\Omega))_{u}$. Moreover, since $u\in\operatorname{dom}A^{\prime}$, the space $\operatorname{dom}A^{\prime}\cap(L^{q}(\Omega))_{u}$ is clearly majorizing in $(L^{q}(\Omega))_{u}$. So it only remains to show that $\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to(L^{q})_{u}}\leq\tilde{c}t^{-\alpha}$ for a constant $\tilde{c}>0$ and all $t\in(0,1]$.
@@
To this end, we first observe that the semigroup $(T(t))_{t\geq 0}$ is dominated by the semigroup generated by the same differential operator but with Neumann boundary conditions; this follows for instance from [50, Corollary 2.22]. In addition to the principal ideals $(L^{p}(\Omega))_{u}$ and $(L^{q}(\Omega))_{u}$ we will now use the weighted $L^{1}$-space $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$. This space coincides with the norm completion of $L^{p}(\Omega)$ with respect to the norm $\left\lVert f\right\rVert_{L^{1}(\mathrm{d}x)}:=\int_{\Omega}\left\lvert f\right\rvert u\;\mathrm{d}x$ and behaves, in a sense, dually to the principal ideal $(L^{q}(\Omega))_{u}$; we refer to [9, Section 2] or [16, Section 2] for a detailed discussion.
@@
Moreover, since $T(t)^{\prime}u=e^{-t\lambda_{0}}u$ for each $t$, the operator $T(t)$ extends to a bounded linear operator on $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$, again denoted by $T(t)$, which satisfies
@@
(an abstract version of this argument can be found in [9, Proposition 2.2]). We now combine the previous two norm estimates with an interpolation inequality: for every function $f\in(L^{p}(\Omega))_{u}\subseteq L^{p}(\Omega)\subseteq L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$, we have
@@
so $\left\lVert f\right\rVert_{L^{p}}\leq\left\lVert u^{p-2}\right\rVert_{\infty}^{1/p}\left\lVert f\right\rVert_{L^{1}(u\;\mathrm{d}x)}^{1/p}\left\lVert f\right\rVert_{(L^{p})_{u}}^{1/q}$; here we have also used that $u\in L^{\infty}(\Omega)$ and $p\geq 2$, which ensures $u^{p-2}\in L^{\infty}(\Omega)$. This estimate for $\left\lVert f\right\rVert_{L^{p}}$ together with the two aforementioned norm estimates for $T(t)$ readily yield
$\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to L^{p}}\leq c^{1/q}\left\lVert u^{p-2}\right\rVert_{\infty}^{1/p}t^{-\alpha}$
for all $t>0$. For every $t>0$, one has $\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to L^{p}}=\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to(L^{q})_{u}}$. Indeed, the inequality $\leq$ is shown in an abstract setting in [16, Proposition 2.1], but one can show that even equality is true by using the duality result in [53, Exercise IV.9(a)]. So the assumption of Theorem 2.12(b) is satisfied, which gives the claimed result.
For Dirichlet boundary conditions, not all elements of $L^{p}(\Omega)_{-1,+}$ can be described as measures on $\overline{\Omega}$. This is in contrast to the situation for Neumann boundary conditions and small $p$, see Example 2.13(b). One can see this by taking $A$ to be the Dirichlet Laplacian on the interval $(-1,1)$. Then $A^{\prime}$ is also the Dirichlet Laplacian and the domain $\operatorname{dom}A^{\prime}$ is (no matter which value we use for $p$) a majorizing and dense subset of $C^{1}[-1,1]$. Hence, $f\mapsto\mp f^{\prime}(\pm 1)$ are positive elements of $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}((-1,1))_{-1}$ that are not given by measures on $[-1,1]$.
3. Admissibility of observation operators
We now come to the main topic of the paper – admissibility. In this section, we are interested in the system
@@ 3. Admissibility of observation operators
where $A$ generates a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on a Banach space $X$ and the observation operator $C\in\mathcal{L}(X_{1},Y)$ takes values in a Banach space $Y$. We say that $C$ is a $\mathrm{L}^{1}$-admissible observation operator if the output map – defined as
$\Psi_{\tau}:X_{1}\to L^{1}([0,\tau],Y),\qquad x\mapsto CT(\mathord{\,\cdot\,})x$ (3.1)
has a bounded extension to $X$ for some (equivalently, all) $\tau>0$. This is of course, equivalent to the existence of $K_{\tau}>0$ such that
$\left\lVert CT(\mathord{\,\cdot\,})x\right\rVert_{L^{1}([0,\tau],Y)}\leq K_{\tau}\left\lVert x\right\rVert\qquad(x\in X_{1}).$ (3.2)
In addition, if we have $\limsup_{\tau\downarrow 0}\left\lVert\Psi_{\tau}\right\rVert_{\mathcal{L}(X,\mathrm{L^{1}}([0,\tau],Y))}=0$, then $C$ is called a zero-class $\mathrm{L^{1}}$-admissible observation operator. Most of our results in this section are proved in the following setting:
Assumption 3.1.
@@
In all of our results, the order on $X_{1}$ is assumed to be the order inherited from $X$. We begin with obtaining sufficient conditions for an operator $C\in\mathcal{L}(X_{1},Y)$ to be $L^{1}$-admissible. Thereafter, in Section 3.2, we look at situations where $L^{1}$-admissibility of $C$ automatically implies zero-class $L^{1}$-admissibility.
3.1. $L^{1}$-admissibility
For positive semigroups on Banach lattices, it is known from [62, Theorem 4.1.17] that positive observation operators mapping into AL-spaces (cf. Table 1) are always $L^{1}$-admissible. In our first result, we generalise this to ordered Banach spaces whose norm is additive on the positive cone. Consequently, we are able to show that every positive finite-rank observation operator is $L^{1}$-admissible (Corollary 3.5).
@@ 3.1. $L^{1}$-admissibility
Suppose that Assumption 3.1 holds. Assume further that $Y$ is an ordered Banach space whose norm is additive on the positive cone and that $C\in\mathcal{L}(X_{1},Y)$ is positive. Then $C$ is $L^{1}$-admissible.
Examples of ordered Banach space whose norm is additive on the positive cone but are not AL-spaces are non-commutative $L^{1}$-spaces, pre-duals of non-commutative von Neumann algebras, or ice-cream cone in finite-dimensions. Some more examples are given in [28, Appendix A].
For the proof of Proposition 3.2, we essentially adapt the argument of [62, Theorem 4.1.17] to our setting. The essential ingredient of the argument is that due to the positivity of the semigroup and the observation operator, it suffices to establish (3.2) for positive $x\in X_{1}$. When $X$ and $Y$ are both Banach lattices, this was proved in [22, Lemma 3.1]. However, the result remains true under our assumption as well and is actually a consequence of the following technical lemma.
@@
Let $X,\widetilde{X}$ be ordered Banach spaces, let the cone $X_{+}$ be generating in $X$, and let $Y$ be a Banach space. Let $(F(t))_{t\in[0,1]}$ be a family of operators in $\mathcal{L}(\widetilde{X},Y)$ with strongly measurable orbits. Let $J\in\mathcal{L}(\widetilde{X},X)_{+}$ and assume the following:
@@
    There exists a sequence $(R_{n})\subseteq\mathcal{L}(X,\widetilde{X})_{+}$ such that $(JR_{n})$ and $(R_{n}J)$ converge strongly to the identity operators on $X$ and $\widetilde{X}$, respectively.
@@
    There exists $\widetilde{M}\geq 0$ such that $\int_{0}^{1}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq\widetilde{M}\left\lVert Jx\right\rVert$ for each $x\in\widetilde{X}_{+}$.
Then there exists $M\geq 0$ such that $\int_{0}^{1}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq M\left\lVert Jx\right\rVert$ even for all $x\in\widetilde{X}$.
Proof.
Let $x\in\widetilde{X}$. By the uniform decomposition property of ordered Banach spaces with generating cones [4, Theorem 2.37], there exists $c>0$ – independent of $x$ – and $y,z\in X_{+}$ such that $Jx=y-z$ and $\left\lVert y\right\rVert,\left\lVert z\right\rVert\leq c\left\lVert Jx\right\rVert$. Due to (b),
$\int_{0}^{1}\left\lVert F(t)R_{n}Jx\right\rVert\;\mathrm{d}t\leq\widetilde{M}\left(\left\lVert JR_{n}y\right\rVert+\left\lVert JR_{n}z\right\rVert\right)$
for all $n\in\mathbb{N}$. Fatou’s lemma and the convergence properties in (a) thus give
$\int_{0}^{1}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq\widetilde{M}\big{(}\left\lVert y\right\rVert+\left\lVert z\right\rVert\big{)}\leq 2c\widetilde{M}\left\lVert Jx\right\rVert,$
as desired. ∎
Proof of Proposition 3.2.
Let $C\in\mathcal{L}(X_{1},Y)$ be positive and assume without loss of generality that $\omega_{0}(A)<0$. Then for every positive $x\in X_{1}$, the integral $\int_{0}^{1}\left\lVert CT(t)x\right\rVert\;\mathrm{d}t$ can be estimated from above by
$\int_{0}^{\infty}\left\lVert CT(t)x\right\rVert\;\mathrm{d}t=\left\lVert\int_{0}^{\infty}CT(t)x\;\mathrm{d}t\right\rVert=\left\lVert CA^{-1}x\right\rVert\leq\left\lVert CA^{-1}\right\rVert\left\lVert x\right\rVert;$
here the first equality holds because the norm on the positive cone of $Y$ is additive. Now we can apply Lemma 3.3 to the space $\tilde{X}:=X_{1}$, the canonical embedding $J\in\mathcal{L}(X_{1},X)_{+}$, the resolvent operators $R_{n}:=nR(n,A)\in\mathcal{L}(X,X_{1})_{+}$, and the operators $F(t):=CT(t)\in\mathcal{L}(X_{1},Y)$. ∎
@@
Suppose that Assumption 3.1 holds. Assume further that there exists an ordered Banach space $\widetilde{X}$ whose norm is additive on the cone $\widetilde{X}_{+}$ such that $C$ factorises as
@@
Proof.
By Proposition 3.2, the operator $C_{1}$ is $L^{1}$-admissible. Thus, for $\tau>0$ there exists $K_{\tau}>0$ such that
@@
for all $x\in X_{1}$, As a result, $C=C_{2}C_{1}$ is also $L^{1}$-admissible. ∎
The factorisation condition in the preceding theorem might seem artificial at first glance. Yet, the next two results show how the condition arises naturally in some situations.
@@
Suppose that Assumption 3.1 holds. If $Y$ is an ordered Banach space and $C\in\mathcal{L}(X_{1},Y)$ is a positive finite-rank operator, then $C$ is $L^{1}$-admissible.
Proof.
It suffices to find an ordered Banach space $\widetilde{X}$ such that the assumptions of Theorem 3.4 hold. Define $\widetilde{X}\subseteq Y$ to be the range $\operatorname{Rg}C$ endowed with the cone $\widetilde{X}_{+}:=\widetilde{X}\cap Y_{+}$. Then $\widetilde{X}$ is a finite-dimensional ordered Banach space and due to finite-dimensionality there exists an equivalent norm on this space that is additive on the positive cone; see [4, Corollary 3.8] and [64, Theorem VII.1.1]. Setting $C_{1}:=C:X_{1}\to\widetilde{X}$ and $C_{2}:\widetilde{X}\to Y$ to be the canonical embedding, Theorem 3.4 can be applied. ∎
We point out that in Corollary 3.5, although $\operatorname{Rg}C$ is a finite-dimensional ordered (Banach) space, it need not be a lattice. So, even though they’re exists an equivalent norm on $\operatorname{Rg}C$ that is additive on the positive cone, there may not be an equivalent norm on $\operatorname{Rg}C$ that turns it into a lattice. For this reason, we needed the generalisation of [62, Theorem 4.1.17] given in Proposition 3.2.
Remark 3.6.
In general, a non-positive finite-rank observation operator need not be $L^{1}$-admissible. Indeed, for the left translation semigroup on $X:=C_{0}[0,1)$, the observation operator $C:X_{1}\to\mathbb{C}$ given by $f\mapsto f^{\prime}(0)$ is not $L^{1}$-admissible.
In fact, even analyticity of the semigroup is not sufficient to guarantee $L^{1}$-admissibility of a finite-rank operator [38, Theorem 10].
The dual of the factorization assumption can be characterised in terms of order boundedness (Appendix A). This lets us reformulate Theorem 3.4 for reflexive $Y$:
@@
Suppose that Assumption 3.1 holds and assume that $Y$ is reflexive. If the dual operator $C^{\prime}:Y^{\prime}\to(X_{1})^{\prime}=(X^{\prime})_{-1}$ maps the unit ball of $Y^{\prime}$ into an order bounded subset of $(X^{\prime})_{-1}$, then $C$ is $L^{1}$-admissible.
Proof.
@@
$C:X_{1}\overset{C_{2}^{\prime}}{\longrightarrow}(\widetilde{X})^{\prime}\overset{C_{1}^{\prime}}{\longrightarrow}Y.$
Finally, the norm on $(\widetilde{X})^{\prime}$ is additive on its positive cone because $\widetilde{X}$ has a unit [48, Lemmata 2 and 3]. Moreover, $C_{2}^{\prime}$ is positive because $C_{2}$ is. In particular, all assumptions of Theorem 3.4 are fulfilled and thus $C$ is $L^{1}$-admissible. ∎
3.2. Zero-class $L^{1}$-admissibility
While $L^{1}$-admissible observation operators need not be zero-class $L^{1}$-admissible (Example 3.11), we are interested in situations when this is indeed the case. To this end, we take advantage of the nilpotency of the right-translation semigroup on $L^{1}([0,1],Y)$ to formulate a simple test for $L^{1}$-admissibility to be zero-class.